# Refereed Learning

### v1

@@ 1.1.3 Our techniques
Next we deal with the case of a general (i.e., not zero-one) metric $\ell$ on the range $\mathcal{Y}$. This case is more challenging since different points in $\mathcal{Y}$ may have very different contributions to the overall loss. In particular, there may be a single point $x^{*}$ with $\ell(h_{1}(x^{*}),f(x^{*}))\gg\ell(h_{0}(x^{*}),f(x^{*}))$, and thus a naive learner which, as in the zero-one case, samples from the set $S={\left\{{x\mid\ell(h_{0}(x),h_{1}(x))>0}\right\}}$ and outputs the hypothesis with better loss on the sample, will require many samples before obtaining the point $x^{*}$ and thus may select the worse model.
To circumvent this issue, we define a rescaled version of $\mathcal{D}$}, denoted ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, that places more mass on points $x$ where $\ell{\big({h_{0}(x),h_{1}(x)}\big)}$ is large. By the triangle inequality, if $\ell{\big({h_{0}(x),h_{1}(x)}\big)}$ is large, then either $\ell{\big({h_{0}(x),f(x)}\big)}$ or $\ell{\big({h_{1}(x),f(x)}\big)}$ must be large as well. Roughly, this technique resolves the above difficulty since if there is such a point $x^{*}$, then the rescaled distribution will place proportionally more mass on $x^{*}$. More generally, we show that under the rescaled distribution ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, with high probability the worse hypothesis accounts for more than half of the combined empirical loss of $h_{0}$ and $h_{1}$ on $f$, computed over only $O{\left({1+\frac{1}{\varepsilon^{2}}}\right)}$ samples. Leveraging the techniques we develop for the zero-one case, we show that the learner, with the help of the provers, can efficiently provide itself with query access to the probability mass function of ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, and thus can execute the certifiable sample protocol to efficiently generate samples from ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$. The full treatment appears in Section 4.2.
@@
We complement the above general-purpose refereed learning protocol, where the provers need exponential time, by demonstrating a refereed learning protocol for a natural learning task where the provers can be efficient. Specifically, we consider the case where $h_{0}$ and $h_{1}$ are promised to be $j$-juntas (i.e., Boolean functions that each depend only on some set of $j\approx\log d$ input coordinates), and the active index sets $J_{0}$ and $J_{1}$ of $h_{0}$ and $h_{1}$ are given as input to all parties. In this setting, we show that the provers can be implemented efficiently. The idea is the following: In the general case, the task that determines the prover runtime is computing the set $S={\left\{{x\mid h_{0}(x)\neq h_{1}(x)}\right\}}$; when $h_{0}$ and $h_{1}$ are promised to be juntas with $j\approx\log d$ active indices, the provers can compute $S$ in time $\mathop{\mathrm{poly}}d$. Since the distribution is uniform, the certifiable sample protocol used in the general case can also be implemented efficiently, and thus the provers can be made to run in time $\mathop{\mathrm{poly}}d$. See Proposition 6.1 for a formal statement and construction of the protocol.
@@
As described earlier, we show several different impossibility results, where each result demonstrates the optimality of a different aspect of the protocols. The first result (Theorem 5.1) demonstrate that without query access to the ground truth $f$, a learner would in general need a prohibitive number of queries. The argument is straightforward: fix $\left\{{h_{0},h_{1}}\right\}$}, sample $b\sim\{0,1\}$ and let $f\leftarrow h_{b}$. Consider the cheating prover that executes the honest protocol, except it “pretends” that $f=h_{1-b}$. As long as the learner does not obtain any sample $x$ with $f(x)\neq h_{1-b}(x)$, it cannot refute the malicious prover’s claim that $h_{1-b}$ has zero loss, and thus cannot determine if it should accept $h_{0}$ or $h_{1}$. The proof that query access to $Q_{\mathcal{D}}$,the PMF of $\mathcal{D}$}, is necessary (Theorem 5.3) follows a similar outline.
Turning to the complexity of the provers, we first observe that, when the provers only have black-box access to $h_{0}$ and $h_{1}$, the provers may need to query $h_{0}$ and $h_{1}$ at $\Omega(2^{d})$ points to find the hypothesis with better loss, up to a multiplicative factor. (Indeed, for all $z\in\{0,1\}^{d}$ let $h_{z}=\mathds{1}{\left[{x=z}\right]}$. Sample hypotheses $z,z^{\prime}\sim\{0,1\}^{d}$ and ground truth $f\sim{\left\{{h_{z},h_{z^{\prime}}}\right\}}$. Any algorithm $\mathcal{A}$} which gets query access to $h_{z},h_{z^{\prime}}$, and $f$ cannot distinguish whether $f=h_{z}$ or $f=h_{z^{\prime}}$ until it queries either $z$ or $z^{\prime}$. Since these are uniformly random points, $\mathcal{A}$} must make $\Omega(2^{d})$ queries.) In Theorem 5.4 we then extend this bound to the case where the provers are given an explicit description of $h_{0}$ and $h_{1}$. This bound proceeds by reduction from 3-SAT: Any refereed learning protocol which guarantees any purely multiplicative bound on the loss of the computed hypothesis can be used to distinguish satisfiable formulas from unsatisfiable ones.
@@ 1.2 Related work
The idea of considering the computational power of a model that involves a weak verifier and two or more provers, one of which is assumed to be honest, goes back to the works of Feige et al. [FST88, FK97], who also observe that the provers can be viewed as competing. Later, Canetti et al. [CRR11, CRR13] and Kol and Raz [KR14] have demonstrated several protocols for delegating arbitrary computations to untrusted servers in that model. These refereed delegation protocols are both simpler and significantly more efficient than ones designed for a single untrusted prover. In fact, one of the protocols in [CRR13], which is sufficiently efficient to be practical, is currently in commercial use within an application where provers are competing strategic agents, and the protocol is used to incentivize the agents to be truthful [AAT+25].
@@
Ergun et al. and later Rothblum et al. [EKR04, RVW13] consider a weak verifier that uses the power of a single untrusted prover to decide whether some huge mathematical object, which is accessible to the verifier only via queries, has some claimed property, or else is far from having the property. Herman and Rothblum [HR22, HR24b, HR24a] consider the case where the object in question is a distribution, and the verifier’s access to the distribution is via obtaining samples. Goldwasser et al. [GRSY21] consider a (single) prover that wishes to convince a suspicious verifier that a given concept has some desirable properties. In all, to the best of our knowledge, this is the first work that considers the power of the two-prover model in the context of learning, testing, or verifying the properties of black-box objects, models being a special case.
A related and very vibrant area of research is that of debate systems where a panel of competing AI agents debate in order to help a weak referee (either a human or another AI agent) obtain a meaningful decision, often with respect to AI safety and alignment. See, e.g., [ICA18, GCW+24]. However, both the methods and the specific goals in these works are very different than the one here.
@@ 2 Framework for refereed learning
Fix score $\mathcal{S}$ and target $\mathcal{T}$} with respect to parameters $k,d\in\mathbb{N}$ and range $\mathcal{Y}$. Let ${\mathcal{H}}\subseteq{\left\{{h:\{0,1\}^{d}\to\mathcal{Y}}\right\}}$ and $\mathds{D}\subseteq{\left\{{{\mathcal{D}}\mid\mathrm{supp}({\mathcal{D}})\subseteq\{0,1\}^{d}}\right\}}$. Fix oracle access models $\mathcal{O}_{0},\mathcal{O}_{1}$ and $\mathcal{O}_{\mathcal{V}}$, slack parameters $\alpha\geq 1$ and $\eta\geq 0$, and soundness error $\beta\geq 0$. A protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} is a $(k,\alpha,\eta,\beta)$-refereed learning protocol (RLP) for $\mathcal{H}$} and $\mathds{D}$ with respect to $\mathcal{S},{\mathcal{T}}$, and oracles $\mathcal{O}_{0},\mathcal{O}_{1}$ and $\mathcal{O}_{\mathcal{V}}$, if for all distributions ${\mathcal{D}}\in\mathds{D}$, functions $f:\{0,1\}^{d}\to\mathcal{Y}$ and $h_{1},\dots,h_{k}\in{\mathcal{H}}$, the following holds:
@@
Definition 2.3 allows the distribution, metric, and functions to take on arbitrary real values. To handle issues with describing and sending arbitrary real-values, we focus on the setting where the distribution $\mathcal{D}$} and the metric $\ell$ are $\lambda$-precise”:
Definition 2.4 (Set $\mathbb{Q}_{\lambda}$, $\lambda$-precise, distribution family $\mathds{D}_{\lambda}$).
Fix $\lambda\in\mathbb{N}$ and define $\mathbb{Q}_{\lambda}\subseteq\mathbb{Q}$ as the set of all rationals $\frac{p}{q}$ such that $\max{\left\{{|p|,|q|}\right\}}\leq 2^{\lambda}$. A metric $\ell$ is $\lambda$-precise if its image satisfies $\mathsf{im}(\ell)\subseteq\mathbb{Q}_{\lambda}$. A probability distribution $\mathcal{D}$} over $\{0,1\}^{d}$ with probability mass function $Q_{\mathcal{D}}$ is $\lambda$-precise if its image satisfies $\mathsf{im}(Q_{\mathcal{D}})\subseteq\mathbb{Q}_{\lambda}$. Let $\mathds{D}_{d,\lambda}$ denote the set of $\lambda$-precise distributions over $\{0,1\}^{d}$. When $d$ is clear from context we will write $\mathds{D}_{\lambda}$ instead of $\mathds{D}_{d,\lambda}$.
Finally, in order to succinctly refer to the set of all functions $\{0,1\}^{d}\to\mathcal{Y}$ and all distributions over $\mathcal{D}$}, we define the following families:
@@ 3.1 Certifiable sample and certifiable sum
Our first tool, Lemma 3.1, allows the verifier to efficiently sample from a distribution that is close to $\mathcal{D}$} given query access to its probability mass function $Q_{\mathcal{D}}$ (defined by $Q_{{\mathcal{D}}}(x)=\operatorname*{{Pr}}_{X\sim{\mathcal{D}}}{\left[{X=x}\right]}$). For all distributions $P$ and $Q$ over domain $\mathcal{X}$, the total-variation distance (TV) distance between $P$ and $Q$ is $d_{\mathrm{TV}}(P,Q)=\sup_{A\subseteq\mathcal{X}}{\left|{P(A)-Q(A)}\right|}$.
@@
There exists a protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} such that for all $d\in\mathbb{N}$, distributions $\mathcal{D}$} over $\{0,1\}^{d}$ with probability mass function $Q_{\mathcal{D}}$, distance $\delta\in(0,1)$, and sample size $m\in\mathbb{N}$, there exists a distribution $\widehat{{\mathcal{D}}}$ with $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}}\big)}\leq\delta$ such that:
@@
    For all $b\in\{0,1\}$ and ${\mathcal{P}}^{*}_{1-b}$, the output ${\left[{{\mathcal{P}}^{Q_{\mathcal{D}}}_{b},{\mathcal{P}}^{*}_{1-b},\mathcal{V}^{Q_{\mathcal{D}}}}\right]}(d,\delta,m)$ consists of $m$ samples $x_{1},\dots,x_{m}\sim\widehat{{\mathcal{D}}}$.
@@
    The verifier’s runtime and protocol’s communication complexity are $(m+\frac{1}{\delta}\log\frac{1}{\delta})\cdot\mathop{\mathrm{poly}}d$.
In order to prove Lemma 3.1 we leverage our second important tool, Lemma 3.2, which gives a protocol for the verifier to efficiently determine the answer to arbitrary functions of the form $\sum_{x\in\{0,1\}^{d}}t(x)$ given only query access to $t$.
@@
Fix $\lambda\in\mathbb{N}$. There exists a protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} such that for all dimensions $d\in\mathbb{N}$ and functions $t:\{0,1\}^{d}\to\mathbb{Q}_{\lambda}$ the following holds:
@@
Let $S=\mathrm{supp}({\mathcal{D}})$ and $N=|S|$. Let $S_{i}$ denote the ${i}^{\mathrm{th}}$ element of $S$ when its elements are ordered according to their probability masses. At a high level, the verifier in the certifiable sample protocol leverages certifiable sum and a protocol called certifiable index (which we present in 3.5) to certifiably construct a monotone distribution $\mathcal{D}$$^{\prime}$ over $[N]$ that captures the “shape” of $\mathcal{D}$}—that is, sampling $i\sim{\mathcal{D}}^{\prime}$ and outputting $x=S_{i}$ yields a distribution $\widehat{{\mathcal{D}}}$ such that $d_{\mathrm{TV}}(\widehat{{\mathcal{D}}},{\mathcal{D}})\leq\delta$. Thus, to sample such an $x$, the verifier simply samples $i\sim{\mathcal{D}}^{\prime}$ and then executes certifiable index to obtain $S_{i}$.
In order to bound the runtime and communication complexity of the verifier, we use a well-known decomposition result from [Bir87] which states that a monotone distribution over $[N]$ can be approximated to error $\delta$ by a histogram with $O(\log(N)/\delta)$ buckets of exponentially increasing cardinality. We present the formulation of Birgé’s decomposition presented in [Can20, Appendix D.4].
Definition 3.3 (Oblivious decomposition, flattened histogram [Can20]).
For all $N\in\mathbb{N}$ and $\delta>0$, the corresponding oblivious decomposition of $[N]$ is the partition $(I_{1},\ldots,I_{\ell})$ of disjoint intervals, where $\ell=\Theta{\left({\frac{\log N}{\delta}}\right)}$, and $|I_{k}|\leq\lfloor(1+\delta)^{k}\rfloor$ for all $k\in[\ell]$.
Additionally, define the flattened histogram $\Phi_{\delta}{\left[{{\mathcal{D}}}\right]}$ as follows:
$\forall k\in[\ell],\forall i\in I_{k},~\Phi_{\delta}{\left[{{\mathcal{D}}}\right]}(i)=\frac{{\mathcal{D}}{\left({I_{k}}\right)}}{|I_{k}|}.$
Fact 3.4 (Birgé’s decomposition [Bir87, Can20]).
If $\mathcal{D}$} is a monotone non-increasing distribution over $[N]$ then $d_{\mathrm{TV}}{\left({{\mathcal{D}},\Phi_{\delta}{\left[{{\mathcal{D}}}\right]}}\right)}\leq\delta$ for all $\delta>0$.
For a set $S$, a total ordering $\prec$, and an algorithm $\mathcal{A}$}, let ${\mathcal{A}}^{S,\prec}$ denote $\mathcal{A}$} with membership query access to $S$ and query access to $\mathds{1}{\left[{\cdot\prec\cdot}\right]}$ (the function that takes as input $(x,x^{\prime})$ and returns $1$ if $x\prec x^{\prime}$ and 0 otherwise).
Claim 3.5 (Certifiable index).
  1. 1.
  2. 2.
@@
We defer the proof of 3.5 and continue with the proof of Lemma 3.1. Since probabilities may require arbitrary precision to express completely, the provers cannot hope to send the exact probabilities. In order to circumvent this issue and bound the communication cost of certifiable sample, we will first perform a preproccessing step that provides the verifier with query access to the probability mass function $Q_{{\mathcal{D}}_{\lambda}}$ of the distribution ${\mathcal{D}}_{\lambda}$ defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$ where $\lambda>d$ is an integer and $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$—that is, $\lfloor{y}\rfloor_{\lambda}$ denotes the nearest multiple of $2^{-\lambda}$ that is at most $y$.
Before presenting the certifiable sample protocol, we introduce an ordering $\prec$ which will ensure the distribution $\mathcal{D}$$^{\prime}$ is monotone so that we can apply Birgé’s decomposition (3.4). Define the ordering $\prec$ on $\{0,1\}^{d}$ by $x\prec y$ if ${\mathcal{D}}_{\lambda}(x)<{\mathcal{D}}_{\lambda}(y)$, with ties broken according to the lexicographical ordering on $\{0,1\}^{d}$.
[…]

### v2

@@ 1.1.3 Our techniques
Next we deal with the case of a general (i.e., not zero-one) metric $\ell$ on the range $\mathcal{Y}$. This case is more challenging since different points in $\mathcal{Y}$ can have very different contributions to the overall loss. In particular, there may be a single point $x^{*}$ with $\ell(h_{1}(x^{*}),f(x^{*}))\gg\ell(h_{0}(x^{*}),f(x^{*}))$, and thus a naive learner which, as in the zero-one case, samples from the set $S={\left\{{x\mid\ell(h_{0}(x),h_{1}(x))>0}\right\}}$ and outputs the hypothesis with better loss on the sample, will require many samples before obtaining the point $x^{*}$ and thus may select the worse model.
To circumvent this issue, we define a rescaled version of $\mathcal{D}$}, denoted ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, that places more mass on points $x$ where $\ell{\big({h_{0}(x),h_{1}(x)}\big)}$ is large. By the triangle inequality, if $\ell{\big({h_{0}(x),h_{1}(x)}\big)}$ is large, then either $\ell{\big({h_{0}(x),f(x)}\big)}$ or $\ell{\big({h_{1}(x),f(x)}\big)}$ must be large as well. Roughly, this technique resolves the above difficulty of finding a higher-loss point $x^{*}$ since if there is such a point, then the rescaled distribution will place proportionally more mass on $x^{*}$. More generally, we show that under the rescaled distribution ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, with high probability the worse hypothesis accounts for more than half of the combined empirical loss of $h_{0}$ and $h_{1}$ on $f$, computed over only $O{\left({1+\frac{1}{\varepsilon^{2}}}\right)}$ samples. Leveraging the techniques we develop for the zero-one case, we show that the learner, with the help of the provers, can efficiently provide itself with query access to the probability mass function of ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$, and thus can execute the certifiable sample protocol to efficiently generate samples from ${\mathcal{D}}^{h_{0},h_{1}}_{\ell}$. The full treatment appears in Section 4.2. Roughly speaking, the statement in Theorem 1.3 holds for the general setting, except we obtain a $(3+\varepsilon,0,\beta)$-refereed learning protocol.
Theorem 1.4 (Refereeed learning for general loss functions (Theorem 4.4, informal)).
There exists a protocol ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ that, for all $\varepsilon,\beta>0$ and distributions $\mathcal{D}$} on $\{0,1\}^{d}$, is a $(3+\varepsilon,0,\beta)$-refereed learning protocol for $f$ and $\mathcal{D}$} with respect to a specified metric $\ell$. The protocol has communication complexity and verifier runtime $\widetilde{O}{\left({(1+\frac{1}{\varepsilon})^{2}\cdot\mathop{\mathrm{poly}}d\cdot\log\frac{1}{\beta}}\right)}$, and the verifier makes 1 query to $f$.
@@
We complement the above general-purpose refereed learning protocols, where the provers need exponential time, by demonstrating a refereed learning protocol for a natural learning task where the provers can be efficient. Specifically, we consider the case where $h_{0}$ and $h_{1}$ are promised to be $j$-juntas (i.e., Boolean functions that each depend only on some set of $j$ input coordinates), and the active index sets $J_{0}$ and $J_{1}$ of $h_{0}$ and $h_{1}$ are given as input to all parties. In this setting, we show that the provers can be implemented efficiently. The idea is the following: In the general case, the task that determines the prover runtime is computing the set $S={\left\{{x\mid h_{0}(x)\neq h_{1}(x)}\right\}}$; when $h_{0}$ and $h_{1}$ are promised to be juntas with $j\approx\log d$ active indices, the provers can compute $S$ in time $\mathop{\mathrm{poly}}d$. Since the distribution is uniform, the certifiable sample protocol used in the general case can also be implemented efficiently, and thus the provers can be made to run in time $\mathop{\mathrm{poly}}d$. See Proposition 6.1 for a formal statement and construction of the protocol.
@@
As described earlier, we show several impossibility results, each demonstrating the optimality of a different aspect of the protocols. The first result (Theorem 5.1) demonstrate that without query access to the ground truth $f$, a learner would in general need a prohibitive number of queries. The argument is straightforward: fix $\left\{{h_{0},h_{1}}\right\}$}, sample $b\sim\{0,1\}$ and let $f\leftarrow h_{b}$. Consider the cheating prover that executes the honest protocol, except it “pretends” that $f=h_{1-b}$. As long as the learner does not obtain any sample $x$ with $f(x)\neq h_{1-b}(x)$, it cannot refute the malicious prover’s claim that $h_{1-b}$ has zero loss, and thus cannot determine if it should accept $h_{0}$ or $h_{1}$. The proof that query access to $Q_{\mathcal{D}}$, the PMF of $\mathcal{D}$}, is necessary (Theorem 5.3) follows a similar outline.
Turning to the complexity of the provers, we first observe that, when the provers only have black-box access to $h_{0}$ and $h_{1}$, the provers may need to query $h_{0}$ and $h_{1}$ at $\Omega(2^{d})$ points to find the hypothesis with better loss, up to a multiplicative factor. (To see this, for all $z\in\{0,1\}^{d}$ let $h_{z}=\mathds{1}{\left[{x=z}\right]}$, and sample hypotheses $z,z^{\prime}\sim\{0,1\}^{d}$ and ground truth $f\sim{\left\{{h_{z},h_{z^{\prime}}}\right\}}$. Any algorithm $\mathcal{A}$} which gets query access to $h_{z},h_{z^{\prime}}$, and $f$ cannot distinguish whether $f=h_{z}$ or $f=h_{z^{\prime}}$ until it queries either $z$ or $z^{\prime}$. Since these are uniformly random points, $\mathcal{A}$} must make $\Omega(2^{d})$ queries.) In Theorem 5.4 we then extend this bound to the case where the provers are given an explicit description of $h_{0}$ and $h_{1}$. This bound proceeds by reduction from 3-SAT: Any refereed learning protocol which guarantees any purely multiplicative bound on the loss of the computed hypothesis can be used to distinguish satisfiable formulas from unsatisfiable ones.
@@ 1.2 Related work
The idea of considering the computational power of a model that involves a weak verifier and two or more provers, one of which is assumed to be honest, goes back to the works of Feige et al. [FeigeST88, FeigeK97], who also observe that the provers can be viewed as competing. [^4] Indeed, with sufficiently large rewards to provers whose proposals are adopted and who demonstrate the other prover’s inconsistencies, as well as harsh penalties to provers who demonstrate inconsistency of the other prover, the provers are incentivized to (a) always provide the best possible model and (b) never provide a demonstrable inconsistency—even without assuming honesty of either prover. Later, [CanettiRR11, CanettiRR13] and [KolR14] have demonstrated several protocols for delegating arbitrary computations to untrusted servers in this model. These refereed delegation protocols are both simpler and significantly more efficient than ones designed for a single untrusted prover. In fact, one of the protocols in [CanettiRR13] is currently in commercial use within an application where provers are competing strategic agents, and the protocol is used to incentivize the agents to be truthful [ArunATQKBEFB25]. Works in the two-server model, such as Prio of [Corrigan-GibbsB17], also often assume honesty from at least one server, though that line of work is closer to threshold MPC than to the refereed computation literature.
@@
The works of [ErgunKR04, RothblumVW13] consider a weak verifier that uses the power of a single untrusted prover to decide whether some huge mathematical object, which is accessible to the verifier only via queries, has some claimed property, or else is far from having the property. [HermanR22, HermanR24-1, HermanR24-2] consider the case where the object in question is a distribution, and the verifier’s access to the distribution is via obtaining samples. [GoldwasserRSY21] consider a (single) prover that wishes to convince a suspicious verifier that a given concept has some desirable properties. In all, to the best of our knowledge, this is the first work that considers the power of the two-prover model in the context of learning, testing, or verifying the properties of black-box objects, models being a special case.
A related and very vibrant area of research is that of debate systems where a panel of competing AI agents debate in order to help a weak referee (either a human or another AI agent) obtain a meaningful decision, often with respect to AI safety and alignment. See, e.g., [IrvingCA18, GuoCWCPCWZ24]. However, both the methods and the specific goals in these works are very different from the ones here.
@@ 2 Framework for refereed learning
Fix score $\mathcal{S}$ and target $\mathcal{T}$} with respect to parameters $k,d\in\mathbb{N}$ and range $\mathcal{Y}$. Let ${\mathcal{H}}\subseteq{\left\{{h:\{0,1\}^{d}\to\mathcal{Y}}\right\}}$ and $\mathds{D}\subseteq{\left\{{{\mathcal{D}}\mid\mathrm{supp}({\mathcal{D}})\subseteq\{0,1\}^{d}}\right\}}$. Fix oracle access models $\mathcal{O}_{0},\mathcal{O}_{1}$ and $\mathcal{O}_{\mathcal{V}}$, slack parameters $\alpha\geq 1$ and $\eta\geq 0$, and soundness error $\beta\geq 0$. A protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} is a $(k,\alpha,\eta,\beta)$-refereed learning protocol (RLP) for $\mathcal{H}$} and $\mathds{D}$ with respect to $\mathcal{S},{\mathcal{T}}$, and oracles $\mathcal{O}_{0},\mathcal{O}_{1}$ and $\mathcal{O}_{\mathcal{V}}$, if for all distributions ${\mathcal{D}}\in\mathds{D}$, (possibly randomized) functions $f:\{0,1\}^{d}\to\mathcal{Y}$ and $h_{1},\dots,h_{k}\in{\mathcal{H}}$, the following holds:
@@
Definition 2.3 allows the distribution, metric, and functions to take on arbitrary real values. To handle issues with describing and sending arbitrary real-values, we focus on the setting where the distribution $\mathcal{D}$} and the metric $\ell$ can be represented exactly with the following discretization. Specifically, where we let $\mathsf{Prec}_{\lambda}:=(-2^{\lambda},2^{\lambda})\cap 2^{-\lambda}\cdot\mathbb{Z}$ denote the set of multiples of $2^{-\lambda}$ bounded in magnitude by $2^{\lambda}$, we say a function $f$ is $\lambda$-precise if its range is $\mathsf{Prec}_{\lambda}$. We say a distribution $\mathcal{D}$} is $\lambda$-precise if its PMF $Q_{\mathcal{D}}$ is $\lambda$-precise. Additionally, let $\mathds{D}_{d,\lambda}$ denote the set of $\lambda$-precise distributions over $\{0,1\}^{d}$ (when $d$ is clear from context we write $\mathds{D}_{\lambda}$ instead of $\mathds{D}_{d,\lambda}$).
Additionally, in order to succinctly refer to the set of all functions $\{0,1\}^{d}\to\mathcal{Y}$ and all distributions over $\mathcal{D}$}, we define the following families:
@@ 3.1 Certifiable sample and certifiable sum
Our first tool, certifiable sample (see Lemma 3.1), allows the verifier to efficiently sample from a distribution that is close to $\mathcal{D}$} given query access to its probability mass function $Q_{\mathcal{D}}$ (defined by $Q_{{\mathcal{D}}}(x)=\operatorname*{{Pr}}_{X\sim{\mathcal{D}}}{\left[{X=x}\right]}$). For all distributions $P$ and $Q$ over domain $\mathcal{X}$, the total-variation distance (TV) distance between $P$ and $Q$ is $d_{\mathrm{TV}}(P,Q)=\sup_{A\subseteq\mathcal{X}}{\left|{P(A)-Q(A)}\right|}$. For a set $S$, a total ordering $\prec$, and an algorithm $\mathcal{A}$}, let ${\mathcal{A}}^{S,\prec}$ denote $\mathcal{A}$} with membership query access to $S$ and query access to $\mathds{1}{\left[{\cdot\prec\cdot}\right]}$ (the function that takes as input $(x,x^{\prime})$ and returns $1$ if $x\prec x^{\prime}$ and 0 otherwise).
@@
Protocol 1 has the property that for all $d\in\mathbb{N}$, distributions $\mathcal{D}$} over $\{0,1\}^{d}$ with probability mass function $Q_{\mathcal{D}}$, distance $\delta\in(0,1)$, and sample size $m\in\mathbb{N}$, there exists a distribution $\widehat{{\mathcal{D}}}$ with $d_{\mathrm{TV}}{\big({\widehat{{\mathcal{D}}},{\mathcal{D}}}\big)}\leq\delta$ such that:
@@
    For all $b\in\{0,1\}$ and ${\mathcal{P}}^{*}_{1-b}$, the output ${\left[{{\mathcal{P}}^{Q_{\mathcal{D}},\prec}_{b},{\mathcal{P}}^{*}_{1-b},\mathcal{V}^{Q_{\mathcal{D}},\prec}}\right]}(d,\delta,m)$ consists of $m$ samples $x_{1},\dots,x_{m}\sim\widehat{{\mathcal{D}}}$.
@@
    The verifier’s runtime and protocol’s communication complexity are $(m\log\frac{1}{\delta})\cdot\mathop{\mathrm{poly}}d$.
In order to prove Lemma 3.1 we leverage our second important tool, certifiable sum, which gives a protocol for the verifier to efficiently determine the answer to arbitrary functions of the form $\sum_{x\in\{0,1\}^{d}}t(x)$ given only query access to $t$.
@@
Fix $\lambda\in\mathbb{N}$. Protocol 2 has the property that for all dimensions $d\in\mathbb{N}$ and functions $t:\{0,1\}^{d}\to\mathsf{Prec}_{\lambda}$ the following holds:
@@
Let $\prec$ denote $\prec_{d}$. At a high level, this algorithm uses the certifiable sum protocol to perform inverse CDF sampling from $\mathcal{D}$}. Roughly, the verifier selects a uniformly random value in $[0,1]$ and then asks for the value of $x$ such that $\sum_{x^{\prime}\prec x}{\mathcal{D}}(x)<p\leq\sum_{x^{\prime}\preceq x}{\mathcal{D}}(x)$. (In reality, to avoid issues related to numerical precision the verifier actually rounds each probability of the PMF to a multiple of $2^{-\lambda}$ and works with a CDF based on the sum of these rounded values.) The verifier uses calls to certifiable sum protocol (Lemma 3.2) to do this entire process efficiently.
@@
Since exactly representing probabilities may require arbitrary precision, the provers cannot hope to send the exact probabilities. In order to circumvent this issue and bound the communication cost of certifiable sample, we will instead obtain $m$ samples from the distribution ${\mathcal{D}}_{\lambda}$ defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$ where $\lambda>d$ is an integer and $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$—that is, $\lfloor{y}\rfloor_{\lambda}$ denotes the nearest multiple of $2^{-\lambda}$ that is at most $y$.
Define the protocol as follows:
    ${\left[{{\mathcal{P}}_{0}^{Q_{\mathcal{D}},\prec},{\mathcal{P}}_{1}^{Q_{\mathcal{D}},\prec},\mathcal{V}^{Q_{\mathcal{D}},\prec}}\right]}(d,\delta,m)$ 1. $\mathcal{V}$: Set $\lambda\leftarrow d+1+\log\frac{1}{\delta}$. Obtain $T_{\lambda}=\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$ via certifiable sum (Lemma 3.2) with $t(x)=\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$. 2. ${\mathcal{P}}_{0},{\mathcal{P}}_{1}$: Let $S\leftarrow\mathrm{supp}({\mathcal{D}})$. 3. Repeat $m$ times: (a) $\mathcal{V}$: Select $p\sim[0,T_{\lambda}]\cap 2^{-\lambda}\cdot\mathbb{Z}$ uniformly at random and send to ${\mathcal{P}}_{0},{\mathcal{P}}_{1}$. (b) For each $b\in\{0,1\}$: i. ${\mathcal{P}}_{b}$: return the largest666according to $\prec$ $\widehat{x}_{b}$ s.t. $\sum_{\begin{subarray}{c}x\in S\text{ s.t.}\\ x\prec\widehat{x}_{b}\end{subarray}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}<p.$ ii. $\mathcal{V}$: Obtain $p^{\mathit{small}}_{b}=\sum_{\begin{subarray}{c}x\in S\text{ s.t.}\\ x\prec\widehat{x}_{b}\end{subarray}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$ via certifiable sum (Lemma 3.2), using $t(x)=\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}\cdot\mathds{1}{\left[{x\in S\wedge x\prec\widehat{x}_{b}}\right]}.$ iii. $\mathcal{V}$: If $p^{\mathit{small}}_{b}<p\leq p^{\mathit{small}}_{b}+\lfloor{{\mathcal{D}}(\widehat{x}_{b})}\rfloor_{\lambda}$ then accept $\widehat{x}_{b}$; otherwise reject. (c) $\mathcal{V}$: Let $\widehat{x}$ denote the first value in $(\widehat{x}_{0},\widehat{x}_{1})$ that was accepted. Output $\widehat{x}$.